%%


\begin{thebibliography}{58}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Allen et~al.(2022)Allen, Iliescu, and Greiff}]{allen2022single}
Mark~S Allen, Dragos Iliescu, and Samuel Greiff. 2022.
\newblock Single item measures in psychological science: A call to action.
\newblock \emph{European journal of psychological assessment}, 38(1):1--5.

\bibitem[{{American Psychiatric Association}(2013)}]{american2013diagnostic}
{American Psychiatric Association}. 2013.
\newblock \emph{Diagnostic and Statistical Manual of Mental Disorders: DSM-5}, volume~5.
\newblock American psychiatric association Washington, DC.

\bibitem[{Antony and Rowa(2005)}]{antony2005evidence}
Martin~M Antony and Karen Rowa. 2005.
\newblock Evidence-based assessment of anxiety disorders in adults.
\newblock \emph{Psychological Assessment}, 17(3):256--266.

\bibitem[{Barlow et~al.(2014)Barlow, Sauer-Zavala, Carl, Bullis, and Ellard}]{barlow2014nature}
David~H Barlow, Shannon Sauer-Zavala, Jenna~R Carl, Jacqueline~R Bullis, and Kristen~K Ellard. 2014.
\newblock The nature, diagnosis, and treatment of neuroticism: Back to the future.
\newblock \emph{Clinical Psychological Science}, 2(3):344--365.

\bibitem[{Breslau(1985)}]{breslau1985depressive}
Naomi Breslau. 1985.
\newblock Depressive symptoms, major depression, and generalized anxiety: A comparison of self-reports on ces-d and results from diagnostic interviews.
\newblock \emph{Psychiatry Research}, 15(3):219--229.

\bibitem[{Brown and Barlow(2014)}]{brown_anxiety_2014}
Timothy~A Brown and David~H Barlow. 2014.
\newblock \emph{Anxiety and {Related} {Disorders} {Interview} {Schedule} for {DSM}-5 ({ADIS}-{5L}): {Lifetime} Version. {Client} Interview Schedule.}
\newblock Oxford University Press.

\bibitem[{Campbell and Fiske(1959)}]{campbell_convergent_1959}
Donald~T Campbell and Donald~W Fiske. 1959.
\newblock Convergent and discriminant validation by the multitrait-multimethod matrix.
\newblock \emph{Psychological Bulletin}, 56(2):81--105.

\bibitem[{Chen et~al.(2000)Chen, Eaton, Gallo, Nestadt, and Crum}]{chen2000empirical}
Li-Shiun Chen, William~W Eaton, Joseph~J Gallo, Gerald Nestadt, and Rosa~M Crum. 2000.
\newblock Empirical examination of current depression categories in a population-based study: symptoms, course, and risk factors.
\newblock \emph{American Journal of Psychiatry}, 157(4):573--580.

\bibitem[{Cohen(1983)}]{cohen1983cost}
Jacob Cohen. 1983.
\newblock The cost of dichotomization.
\newblock \emph{Applied Psychological Measurement}, 7(3):249--253.

\bibitem[{Connor et~al.(2000)Connor, Davidson, Churchill, Sherwood, Weisler, and Foa}]{connor2000psychometric}
Kathryn~M Connor, Jonathan~RT Davidson, L~Erik Churchill, Andrew Sherwood, Richard~H Weisler, and Edna Foa. 2000.
\newblock Psychometric properties of the social phobia inventory ({SPIN}): New self-rating scale.
\newblock \emph{The British Journal of Psychiatry}, 176(4):379--386.

\bibitem[{Coyne(1994)}]{coyne1994self}
James~C Coyne. 1994.
\newblock Self-reported distress: Analog or ersatz depression?
\newblock \emph{Psychological Bulletin}, 116(1):29--45.

\bibitem[{Creed et~al.(2016)Creed, Wolk, Feinberg, Evans, and Beck}]{creed2016beyond}
Torrey~A Creed, Courtney~Benjamin Wolk, Betsy Feinberg, Arthur~C Evans, and Aaron~T Beck. 2016.
\newblock Beyond the label: Relationship between community therapists’ self-report of a cognitive behavioral therapy orientation and observed skills.
\newblock \emph{Administration and Policy in Mental Health and Mental Health Services Research}, 43:36--43.

\bibitem[{Cronbach(1951)}]{cronbach1951coefficient}
Lee~J Cronbach. 1951.
\newblock Coefficient alpha and the internal structure of tests.
\newblock \emph{Psychometrika}, 16(3):297--334.

\bibitem[{Cronbach and Meehl(1955)}]{cronbach_construct_1955}
Lee~J Cronbach and Paul~E. Meehl. 1955.
\newblock Construct validity in psychological tests.
\newblock \emph{Psychological Bulletin}, 52(4):281--302.

\bibitem[{Cuijpers et~al.(2004)Cuijpers, de~Graaf, and van Dorsselaer}]{cuijpers2004minor}
Pim Cuijpers, Ron de~Graaf, and Saskia van Dorsselaer. 2004.
\newblock Minor depression: risk profiles, functional disability, health care use and risk of developing major depression.
\newblock \emph{Journal of Affective Disorders}, 79(1-3):71--79.

\bibitem[{DeYoung et~al.(2022)DeYoung, Chmielewski, Clark, Condon, Kotov, Krueger, Lynam, Markon, Miller, Mullins-Sweatt et~al.}]{deyoung2022distinction}
Colin~G DeYoung, Michael Chmielewski, Lee~Anna Clark, David~M Condon, Roman Kotov, Robert~F Krueger, Donald~R Lynam, Kristian~E Markon, Joshua~D Miller, Stephanie~N Mullins-Sweatt, et~al. 2022.
\newblock The distinction between symptoms and traits in the hierarchical taxonomy of psychopathology (hitop).
\newblock \emph{Journal of Personality}, 90(1):20--33.

\bibitem[{Eddy et~al.(2008)Eddy, Dorer, Franko, Tahilani, Thompson-Brenner, and Herzog}]{eddy2008diagnostic}
Kamryn~T Eddy, David~J Dorer, Debra~L Franko, Kavita Tahilani, Heather Thompson-Brenner, and David~B Herzog. 2008.
\newblock Diagnostic crossover in anorexia nervosa and bulimia nervosa: Implications for {DSM-V}.
\newblock \emph{American Journal of Psychiatry}, 165(2):245--250.

\bibitem[{Ernala et~al.(2019)Ernala, Birnbaum, Candan, Rizvi, Sterling, Kane, and De~Choudhury}]{ernala2019methodological}
Sindhu~Kiranmai Ernala, Michael~L Birnbaum, Kristin~A Candan, Asra~F Rizvi, William~A Sterling, John~M Kane, and Munmun De~Choudhury. 2019.
\newblock Methodological gaps in predicting mental health states from social media: Triangulating diagnostic signals.
\newblock In \emph{Proceedings of the 2019 {CHI} Conference on Human Factors in Computing Systems}, pages 1--16.

\bibitem[{First et~al.(2016)First, Williams, Karg, and Spitzer}]{first_scid_2016}
Michael~B. First, Janet B.~W. Williams, Rhonda~S. Karg, and Robert~L. Spitzer. 2016.
\newblock \emph{{SCID}-5-{CV}: {Structured} {Clinical} {Interview} for {DSM}-5 {Disorders}: {Clinician} {Version}}.
\newblock American Psychiatric Association Publishing, Arlington, VA.

\bibitem[{Fried(2017{\natexlab{a}})}]{fried2017dep:heterogeneity}
Eiko~I Fried. 2017{\natexlab{a}}.
\newblock Moving forward: How depression heterogeneity hinders progress in treatment and research.
\newblock \emph{Expert Review of Neurotherapeutics}, 17(5):423--425.

\bibitem[{Fried(2017{\natexlab{b}})}]{fried_what_2017}
Eiko~I. Fried. 2017{\natexlab{b}}.
\newblock \href {https://doi.org/10.1080/17437199.2017.1306718} {What are psychological constructs? {On} the nature and statistical modelling of emotions, intelligence, personality traits and mental disorders}.
\newblock \emph{Health Psychology Review}, 11(2):130--134.

\bibitem[{Fried and Nesse(2015)}]{fried2015depression}
Eiko~I Fried and Randolph~M Nesse. 2015.
\newblock Depression is not a consistent syndrome: An investigation of unique symptom patterns in the {STAR*D} study.
\newblock \emph{Journal of Affective Disorders}, 172:96--102.

\bibitem[{Guntuku et~al.(2017)Guntuku, Yaden, Kern, Ungar, and Eichstaedt}]{guntuku_detecting_2017}
Sharath~Chandra Guntuku, David~B Yaden, Margaret~L Kern, Lyle~H Ungar, and Johannes~C Eichstaedt. 2017.
\newblock \href {https://doi.org/10.1016/j.cobeha.2017.07.005} {Detecting depression and mental illness on social media: {An} integrative review}.
\newblock \emph{Current Opinion in Behavioral Sciences}, 18:43--49.

\bibitem[{Hallgren(2012)}]{hallgren2012computing}
Kevin~A Hallgren. 2012.
\newblock Computing inter-rater reliability for observational data: an overview and tutorial.
\newblock \emph{Tutorials in Quantitative Methods for Psychology}, 8(1):23.

\bibitem[{Harvey et~al.(2004)Harvey, Watkins, and Mansell}]{harvey2004cognitive}
Allison~G Harvey, Edward Watkins, and Warren Mansell. 2004.
\newblock \emph{Cognitive behavioural processes across psychological disorders: A transdiagnostic approach to research and treatment}.
\newblock Oxford University Press, USA.

\bibitem[{Haslam et~al.(2012)Haslam, Holland, and Kuppens}]{haslam_categories_2012}
N.~Haslam, E.~Holland, and P.~Kuppens. 2012.
\newblock \href {https://doi.org/10.1017/S0033291711001966} {Categories versus dimensions in personality and psychopathology: {A} quantitative review of taxometric research}.
\newblock \emph{Psychological Medicine}, 42(5):903--920.

\bibitem[{Hunt et~al.(2003)Hunt, Auriemma, and Cashaw}]{hunt2003self}
Melissa Hunt, Joseph Auriemma, and Ashara~CA Cashaw. 2003.
\newblock Self-report bias and underreporting of depression on the {BDI-II}.
\newblock \emph{Journal of Personality Assessment}, 80(1):26--30.

\bibitem[{Inkster et~al.(2016)Inkster, Stillwell, Kosinski, and Jones}]{inkster2016decade}
Becky Inkster, David Stillwell, Michal Kosinski, and Peter Jones. 2016.
\newblock A decade into facebook: Where is psychiatry in the digital age?
\newblock \emph{The Lancet Psychiatry}, 3(11):1087--1090.

\bibitem[{Inoue et~al.(2012)Inoue, Tanaka, Nakagawa, Nakato, Kameyama, Boku, Toda, Kurita, and Koyama}]{inoue2012utility}
Takeshi Inoue, Teruaki Tanaka, Shin Nakagawa, Yasuya Nakato, Rie Kameyama, Shuken Boku, Hiroyuki Toda, Tsugiko Kurita, and Tsukasa Koyama. 2012.
\newblock Utility and limitations of {PHQ-9} in a clinic specializing in psychiatric care.
\newblock \emph{BMC Psychiatry}, 12:1--6.

\bibitem[{Insel et~al.(2010)Insel, Cuthbert, Garvey, Heinssen, Pine, Quinn, Sanislow, and Wang}]{insel2010research}
Thomas Insel, Bruce Cuthbert, Marjorie Garvey, Robert Heinssen, Daniel~S Pine, Kevin Quinn, Charles Sanislow, and Philip Wang. 2010.
\newblock Research domain criteria ({RDoC}): Toward a new classification framework for research on mental disorders.
\newblock \emph{American Journal of Psychiatry}, 167(7):748--751.

\bibitem[{Joiner et~al.(2025)Joiner, Robison, Manwaring, Rienecke, Le~Grange, Duffy, Mehler, and Blalock}]{joiner2025reliability}
Thomas~E Joiner, Morgan Robison, Jamie Manwaring, Renee~D Rienecke, Daniel Le~Grange, Alan Duffy, Philip~S Mehler, and Dan~V Blalock. 2025.
\newblock The reliability and construct validity of a single-item assessment of suicidal ideation among patients with mood and anxiety disorders.
\newblock \emph{Journal of Clinical Psychology}.
\newblock Advanced online publication.

\bibitem[{Judd et~al.(1997)Judd, Akiskal, and Paulus}]{judd1997role}
Lewis~L Judd, Hagop~S Akiskal, and Martin~P Paulus. 1997.
\newblock The role and clinical significance of subsyndromal depressive symptoms ({SSD}) in unipolar major depressive disorder.
\newblock \emph{Journal of Affective Disorders}, 45(1-2):5--18.

\bibitem[{Kendall et~al.(1987)Kendall, Hollon, Beck, Hammen, and Ingram}]{kendall1987issues}
Philip~C Kendall, Steven~D Hollon, Aaron~T Beck, Constance~L Hammen, and Rick~E Ingram. 1987.
\newblock Issues and recommendations regarding use of the {Beck Depression Inventory}.
\newblock \emph{Cognitive Therapy and Research}, 11:289--299.

\bibitem[{Kessler(1997)}]{kessler-1997-comorbidity}
Ronald~C. Kessler. 1997.
\newblock The prevalence of psychiatric comorbidity.
\newblock In Scott~Ed Wetzler and William~C Sanderson, editors, \emph{Treatment strategies for patients with psychiatric comorbidity.}, pages 23--48. John Wiley \& Sons Inc.

\bibitem[{Kessler et~al.(2006)Kessler, Chiu, Jin, Ruscio, Shear, and Walters}]{kessler_epidemiology_2006}
Ronald~C. Kessler, Wai~Tat Chiu, Robert Jin, Ayelet~Meron Ruscio, Katherine Shear, and Ellen~E. Walters. 2006.
\newblock \href {https://doi.org/10.1001/archpsyc.63.4.415} {The epidemiology of panic attacks, panic disorder, and agoraphobia in the {National} {Comorbidity} {Survey} {Replication}}.
\newblock \emph{Archives of General Psychiatry}, 63(4):415--424.

\bibitem[{Klein et~al.(2005)Klein, Dougherty, and Olino}]{klein2005toward}
Daniel~N Klein, Lea~R Dougherty, and Thomas~M Olino. 2005.
\newblock Toward guidelines for evidence-based assessment of depression in children and adolescents.
\newblock \emph{Journal of Clinical Child and Adolescent Psychology}, 34(3):412--432.

\bibitem[{Kotov et~al.(2015)Kotov, Perlman, Gámez, and Watson}]{Kotov_Perlman_Gámez_Watson_2015}
R.~Kotov, G.~Perlman, W.~Gámez, and D.~Watson. 2015.
\newblock \href {https://doi.org/10.1017/S0033291714002815} {The structure and short-term stability of the emotional disorders: a dimensional approach}.
\newblock \emph{Psychological Medicine}, 45(8):1687–1698.

\bibitem[{Kotov et~al.(2017)Kotov, Krueger, Watson, Achenbach, Althoff, Bagby, Brown, Carpenter, Caspi, Clark et~al.}]{kotov2017hierarchical}
Roman Kotov, Robert~F Krueger, David Watson, Thomas~M Achenbach, Robert~R Althoff, R~Michael Bagby, Timothy~A Brown, William~T Carpenter, Avshalom Caspi, Lee~Anna Clark, et~al. 2017.
\newblock The hierarchical taxonomy of psychopathology (hitop): A dimensional alternative to traditional nosologies.
\newblock \emph{Journal of Abnormal Psychology}, 126(4):454.

\bibitem[{Kroenke et~al.(2001)Kroenke, Spitzer, and Williams}]{kroenke1999phq9}
K.~Kroenke, R.~L. Spitzer, and J.~B. Williams. 2001.
\newblock The phq-9: validity of a brief depression severity measure.
\newblock \emph{Journal of General Internal Medicine}, 16:606--613.

\bibitem[{Linn(1968)}]{linn1968range}
Robert~L Linn. 1968.
\newblock Range restriction problems in the use of self-selected groups for test validation.
\newblock \emph{Psychological Bulletin}, 69(1):69--73.

\bibitem[{Markon et~al.(2011)Markon, Chmielewski, and Miller}]{markon2011reliability}
Kristian~E Markon, Michael Chmielewski, and Christopher~J Miller. 2011.
\newblock The reliability and validity of discrete and continuous measures of psychopathology: a quantitative review.
\newblock \emph{Psychological Bulletin}, 137(5):856--879.

\bibitem[{Nook et~al.(2022)Nook, Hull, Nock, and Somerville}]{nook2022linguistic}
Erik~C Nook, Thomas~D Hull, Matthew~K Nock, and Leah~H Somerville. 2022.
\newblock Linguistic measures of psychological distance track symptom levels and treatment outcomes in a large set of psychotherapy transcripts.
\newblock \emph{Proceedings of the National Academy of Sciences}, 119(13):e2114737119.

\bibitem[{Parry et~al.(2010)Parry, Castonguay, Borkovec, and Wolf}]{parry2010practice}
Glenys Parry, Louis~G Castonguay, Tom~D Borkovec, and Abraham~W Wolf. 2010.
\newblock Practice research networks and psychological services research in the uk and usa.
\newblock \emph{Developing and delivering practice-based evidence: A guide for the psychological therapies}, pages 311--325.

\bibitem[{Reichelt et~al.(2003)Reichelt, James, and Blackburn}]{reichelt2003impact}
F~Katharina Reichelt, Ian~A James, and Ivy-Marie Blackburn. 2003.
\newblock Impact of training on rating competence in cognitive therapy.
\newblock \emph{Journal of Behavior Therapy and Experimental Psychiatry}, 34(2):87--99.

\bibitem[{Ruscio(2019)}]{ruscio2019normal}
Ayelet~Meron Ruscio. 2019.
\newblock Normal versus pathological mood: Implications for diagnosis.
\newblock \emph{Annual Review of Clinical Psychology}, 15(1):179--205.

\bibitem[{Ruscio et~al.(2008)Ruscio, Brown, Chiu, Sareen, and Kessler}]{ruscio_social_2008}
Ayelet~Meron Ruscio, Timothy~A Brown, Wai~Tat Chiu, Jitender Sareen, and Ronald~C Kessler. 2008.
\newblock \href {https://doi.org/10.1017/s0033291707001699} {Social fears and social phobia in the {United} {States}: {Results} from the {National} {Comorbidity} {Survey} {Replication}}.
\newblock \emph{Psychological Medicine}, 38(1):15--28.

\bibitem[{Ruscio et~al.(2007)Ruscio, Chiu, Roy-Byrne, Stang, Stein, Wittchen, and Kessler}]{ruscio_broadening_2007}
Ayelet~Meron Ruscio, Wai~Tat Chiu, Peter Roy-Byrne, Paul~E. Stang, Dan~J. Stein, Hans-Ulrich Wittchen, and Ronald~C. Kessler. 2007.
\newblock \href {https://doi.org/10.1016/j.janxdis.2006.10.004} {Broadening the definition of generalized anxiety disorder: {Effects} on prevalence and associations with other disorders in the {National} {Comorbidity} {Survey} {Replication}}.
\newblock \emph{Journal of Anxiety Disorders}, 21(5):662--676.

\bibitem[{Ruscio and Ruscio(2000)}]{ruscio2000informing}
John Ruscio and Ayelet~Meron Ruscio. 2000.
\newblock Informing the continuity controversy: A taxometric analysis of depression.
\newblock \emph{Journal of Abnormal Psychology}, 109(3):473--487.

\bibitem[{Sanchez-Villegas et~al.(2008)Sanchez-Villegas, Schlatter, Ortuno, Lahortiga, Pla, Benito, and Martinez-Gonzalez}]{sanchez2008validity}
Almudena Sanchez-Villegas, Javier Schlatter, Felipe Ortuno, Francisca Lahortiga, Jorge Pla, Silvia Benito, and Miguel~A Martinez-Gonzalez. 2008.
\newblock Validity of a self-reported diagnosis of depression among participants in a cohort study using the {Structured Clinical Interview} for {DSM-IV} ({SCID-I}).
\newblock \emph{BMC Psychiatry}, 8:1--8.

\bibitem[{Shear and Maser(1994)}]{shear1994standardized}
M~Katherine Shear and Jack~D Maser. 1994.
\newblock Standardized assessment for panic disorder research: A conference report.
\newblock \emph{Archives of General Psychiatry}, 51(5):346--354.

\bibitem[{Spitzer et~al.(2006)Spitzer, Kroenke, Williams, and L{\"o}we}]{spitzer2006brief}
Robert~L Spitzer, Kurt Kroenke, Janet~BW Williams, and Bernd L{\"o}we. 2006.
\newblock A brief measure for assessing generalized anxiety disorder: The {GAD-7}.
\newblock \emph{Archives of Internal Medicine}, 166(10):1092--1097.

\bibitem[{Stade et~al.(2023{\natexlab{a}})Stade, DeRubeis, Ungar, and Ruscio}]{stade2023transdiagnostic}
Elizabeth~C Stade, Robert~J DeRubeis, Lyle Ungar, and Ayelet~Meron Ruscio. 2023{\natexlab{a}}.
\newblock A transdiagnostic, dimensional classification of anxiety shows improved parsimony and predictive noninferiority to {DSM}.
\newblock \emph{Journal of Psychopathology and Clinical Science}, 132(8):937–948.

\bibitem[{Stade et~al.(2023{\natexlab{b}})Stade, Ungar, Eichstaedt, Sherman, and Ruscio}]{stade2023depression}
Elizabeth~C Stade, Lyle Ungar, Johannes~C Eichstaedt, Garrick Sherman, and Ayelet~Meron Ruscio. 2023{\natexlab{b}}.
\newblock Depression and anxiety have distinct and overlapping language patterns: Results from a clinical interview.
\newblock \emph{Journal of Psychopathology and Clinical Science}, 132(8):972--983.

\bibitem[{Stanton et~al.(2020)Stanton, McDonnell, Hayden, and Watson}]{stanton2020transdiagnostic}
Kasey Stanton, Christina~G McDonnell, Elizabeth~P Hayden, and David Watson. 2020.
\newblock Transdiagnostic approaches to psychopathology measurement: Recommendations for measure selection, data analysis, and participant recruitment.
\newblock \emph{Journal of Abnormal Psychology}, 129(1):21.

\bibitem[{Stuart et~al.(2014)Stuart, Pasco, Jacka, Brennan, Berk, and Williams}]{stuart2014comparison}
Amanda~L Stuart, Julie~A Pasco, Felice~N Jacka, Sharon~L Brennan, Michael Berk, and Lana~J Williams. 2014.
\newblock Comparison of self-report and structured clinical interview in the identification of depression.
\newblock \emph{Comprehensive Psychiatry}, 55(4):866--869.

\bibitem[{Watson et~al.(2012)Watson, O’Hara, Naragon-Gainey, Koffel, Chmielewski, Kotov, Stasik, and Ruggero}]{watson2012development}
David Watson, Michael~W O’Hara, Kristin Naragon-Gainey, Erin Koffel, Michael Chmielewski, Roman Kotov, Sara~M Stasik, and Camilo~J Ruggero. 2012.
\newblock Development and validation of new anxiety and bipolar symptom scales for an expanded version of the {IDAS} (the {IDAS-II}).
\newblock \emph{Assessment}, 19(4):399--420.

\bibitem[{Watson et~al.(1995)Watson, Weber, Assenheimer, Clark, Strauss, and McCormick}]{watson1995testing}
David Watson, Kris Weber, Jana~Smith Assenheimer, Lee~Anna Clark, Milton~E Strauss, and Richard~A McCormick. 1995.
\newblock Testing a tripartite model: I. evaluating the convergent and discriminant validity of anxiety and depression symptom scales.
\newblock \emph{Journal of Abnormal Psychology}, 104(1):3--14.

\bibitem[{{World Health Organization}(2004)}]{whoicd10}
{World Health Organization}. 2004.
\newblock \href {https://iris.who.int/handle/10665/42980} {\emph{ICD-10 : International Statistical Classification of Diseases and Related Health Problems: Tenth Revision, 2nd Edition}}.
\newblock World Health Organization.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}

\usepackage{geometry}
\usepackage{array}
\usepackage{booktabs}
\usepackage{pdflscape}
\newcommand{\xhdr}[1]{\vspace{1mm}\noindent{{\bf #1.}}}
\usepackage[utf8]{inputenc}

\newcommand{\remove}[1]{}

\title{Measuring Mental Health Variables in Computational Research:\\Toward Validated, Dimensional, and Transdiagnostic Approaches}

\author{Chen Shani\\
  Computer Science\\
  Stanford University\\
  \href{mailto:cshani@stanford.edu}{cshani@stanford.edu} 
  \And
  Elizabeth C. Stade \\
  Human-Centered Artificial Intelligence\\
  Stanford University\\
  \href{mailto:ecs@stanford.edu}{ecs@stanford.edu}
  }
\begin{document}
\maketitle
\begin{abstract}

Computational mental health research develops models to predict and understand psychological phenomena, but often relies on inappropriate measures of psychopathology constructs, undermining validity. We identify three key issues: (1) reliance on unvalidated measures (e.g., self-declared diagnosis) over validated ones (e.g., diagnosis by clinician); (2) treating mental health constructs as categorical rather than dimensional; and (3) focusing on disorder-specific constructs instead of transdiagnostic ones. We outline the benefits of using validated, dimensional, and transdiagnostic measures and offer practical recommendations for practitioners. Using valid measures that reflect the nature and structure of psychopathology is essential for computational mental health research.
\end{abstract}

\section{Introduction}
In computational mental health research, significant effort is invested in designing models to predict and understand psychological phenomena. Yet the validity and utility of these models can be undermined when they rely on flawed or inappropriate representations of psychopathology (i.e., mental disorder). For example, a classifier cannot validly predict depression if its training data is based on an invalid measure of depression.

Measuring psychopathology \textit{constructs} -- representations of psychological states or processes like ``depression'' or ``neuroticism'' \cite{fried_what_2017} -- is challenging. Constructs are abstract and cannot be directly observed (e.g., there is no single biological indicator for depression) and diagnostic systems like \textit{DSM-5} \cite{american2013diagnostic} and \textit{ICD-10} \cite{whoicd10} have slight variations in how they define syndromes. 

Yet clinical psychology has for decades invested in validating  measures of psychopathology \cite{cronbach_construct_1955} and advancing measurement techniques that reflect developments in psychological science \cite{stanton2020transdiagnostic}. Recommended measurement techniques include clinician-administered interviews, self-report questionnaires, and informant reports \cite{stanton2020transdiagnostic}. While no measurement scheme is without error, techniques exist to help ensure that a measure actually taps the construct it purports to (a process known as \textit{construct validity} (\citet{campbell_convergent_1959}).


However, computational psychopathology research has slow to adopt measurement techniques from clinical psychological science, hindering progress. Here, we highlight three key concepts from clinical science—measurement validity, dimensional measurement, and transdiagnostic measurement. Acknowledging that
computational research often turns to inappropriate measures of psychopathology constructs due to the constraints of computational research -- collecting large-scale clinician-assessed data is expensive and time-consuming, and existing or archival data can be difficult or impossible to access due to privacy concerns -- we offer practical recommendations for improved assessment of psychological phenomena in computational research.

\section{Measurement Validity}
Computational research often infers mental health conditions using methods with poor or unknown validity. For example, some studies assume a diagnosis based on forum membership (e.g., \texttt{r/Depression}) or self-declarations (e.g., ``I have depression'') on social media \cite{guntuku_detecting_2017}, ``proxy diagnostic signals'' which have been shown to have poor external validity \cite{ernala2019methodological}. In other cases, computational researchers write their own single-item measure of a psychological construct, rather than selecting an existing measure with good validity \cite{allen2022single}. The clinical science literature provides justification that all measurement methods are not equal vis-a-vis validity: For instance, self-reported history of depression diagnosis only modestly agrees with semi-structured diagnostic interview findings \cite{stuart2014comparison, sanchez2008validity}. 


To mitigate this, computational researchers should use measures of psychological constructs with good validity. Gold standard psychopathology assessment typically involves a clinician-administered structured or semi-structured interview (e.g., Structured Clinical Interview for \textit{DSM-5}; \citet{first_scid_2016} based on an established psychopathology classification system (e.g., \textit{DSM-5}. As an alternative, self- or informant-report measures that have undergone rigorous psychometric evaluation, such as the PHQ-9 \cite{kroenke1999phq9}, can be used \cite{stanton2020transdiagnostic}.

\subsection{Limitations of Self-Report Questionnaires}

Self-report methodology, while offering high levels of convenience,  has meaningful limitations of which computational researchers should be aware:

\xhdr{Bias} Self-report measures are vulnerable to biased patterns of responding such as participant lack of insight or yay- or nay-saying biases \cite{hunt2003self}, which can introduce systematic errors into computational models.

\xhdr{Specificity of Constructs} Self-report tools may be imprecise measures of psychological constructs. Evidence exists that putative self-report measures of depression may in fact capture general distress \cite{coyne1994self, kendall1987issues} or anxiety \cite{breslau1985depressive} rather than just depression. This lack of precision can weaken model predictions and blur construct boundaries.

\xhdr{Using Tools Outside Their Intended Setting} Self-report measures may be less effective when used outside their original context. For example, the PHQ-9, designed for primary care, has low specificity and PPV in specialty mental health settings \cite{inoue2012utility}, limiting its validity in clinical samples. Developed as a screening tool, it maps perfectly onto \textit{DSM-5} criteria, but it does not assess other symptoms associated with depression, like self-dislike or low libido, meaning it has a narrower possible range than measures that capture a wide range of symptoms. Using tools beyond their intended purpose may reduce model accuracy, and range restriction can attenuate effects (see below).


\section{Dimensional Measurement}

Mental health constructs can be assessed dimensionally (e.g., on a scale from 0 [\textit{no depression}] to 10 [\textit{extreme depression}]) or using categorical labels (\textit{not depressed} vs. \textit{depressed}). Despite computational research tending to employ categorical measurement schemes, most psychopathology constructs are inherently dimensional, as evidenced by the following:

At the \textbf{manifest (observable) level}, symptoms show a monotonic relationship with functional outcomes (\citet{kessler_epidemiology_2006, ruscio_broadening_2007, ruscio_social_2008, cuijpers2004minor, judd1997role}). Even mild or infrequent depression symptoms that fall below the \textit{DSM-5} criteria for major depressive disorder are associated with impairment. However, categorical representations of psychopathology group all subthreshold symptom presentations together, obscuring mild yet clinically meaningful dysfunction \cite{ruscio2019normal}.

Furthermore, longitudinal research reveals that individuals frequently fluctuate between levels of severity of symptoms over time, including crossing thresholds above and below thresholds week-by-week \cite{chen2000empirical, judd1997role}, making diagnoses somewhat arbitrary depending on the time of evaluation.

At the \textbf{latent level}, taxometric analysis of psychometric variables, which compares categorical and dimensional models \cite{ruscio2000informing}, typically yields dimensional solutions. This indicates that constructs like depression and anxiety have a natural, underlying structure that is dimensional, not categorical \cite{haslam_categories_2012}.

To reflect the true nature and structure of psychopathology, computational researchers should treat most mental health variables as continuous. Using dimensional measures of psychopathology can improve granularity of computational models (e.g., allowing models to differentiate between moderate and severe symptom presentations), improving accuracy and clinical utility.

\section{Transdiagnostic Measurement}

While diagnostic systems like the \textit{DSM-5} dominate clinical practice and research, there is growing recognition of \textit{transdiagnostic processes} that cut across the mental disorders \citep{harvey2004cognitive}. For example, \textit{avoidance} is shared by many different types of anxiety disorders, including panic disorder, specific phobia, and social anxiety disorder. Growing evidence supports transdiagnostic conceptualizations of psychopathology, including:

\xhdr{High Comorbidity} Mental disorders frequently co-occur at rates far exceeding chance. For example, people diagnosed with one anxiety disorder are six times more likely to have another \cite{kessler-1997-comorbidity}, suggesting shared underlying mechanisms.

\xhdr{High Diagnostic Crossover} Many patients transition between diagnoses over time. For example, 20-50\% of individuals diagnosed with anorexia nervosa later develop bulimia \cite{eddy2008diagnostic}, suggesting that existing diagnostic categories may sub-optimally represent psychopathology.

\xhdr{Non-Specific Treatment Effects} Treatments targeting one disorder often alleviate symptoms of co-occurring conditions—for example, PTSD treatments frequently reduce depression symptoms \cite{barlow2014nature}. This suggests that interventions may be acting on transdiagnostic mechanisms rather than disorder-specific factors.

In response, the field has introduced transdiagnostic classification models, such as \textbf{NIMH's Research Domain Criteria (RDoC)} \cite{insel2010research}, which defines cross-cutting dimensions like ``reward responsiveness'' and "potential threat," and the \textbf{Hierarchical Taxonomy of Psychopathology (HiTOP)} \cite{kotov2017hierarchical}, which groups together frequently co-occurring symptoms. 



Transdiagnostic measures offer increased parsimony and can better reflect the psychopathology vis-\`a-vis syndrome-specific alternatives (e.g., \citet{stade2023transdiagnostic}; See \citet{stanton2020transdiagnostic} for guidance on selecting transdiagnostic measures).

\section{Impact of Poor Measurement}
Using measures with poor validity, or using categorical or syndrome-specific measures, can negatively impact computational research. Key implications include:

\xhdr{Reduced Resolution} Using categorical labels for psychopathology oversimplifies complex constructs, discarding valuable information about symptom severity, which is important for model accuracy and has clinical utility (e.g., the difference between moderate and severe depression is meaningful to clinicians).

\xhdr{Mis-Classifying Boundary Cases} Relatedly, categorical representation of psychopathology risks misclassifying individuals who fall close to the diagnostic boundary. 

\xhdr{Risk of Type II Errors} Dichotomizing variables that are continuous in nature sacrifices statistical power \cite{cohen1983cost} and reduces reliability \citep{markon2011reliability}, increasing the risk of Type II errors.

\xhdr{Overfitting and Poor Generalization} Noisy measurements cause models to learn spurious patterns, reducing reliability and real-world applicability.

\xhdr{Misleading Interpretations} Poor measurements can cause misleading conclusions about mental health constructs, such as conflating depression overlapping yet different constructs, like anxiety or negative emotionality.

\xhdr{Erosion of Clinical Utility} For computational models to have practical relevance in mental health care, they must provide insights or predictions that clinicians can act upon. Models based on bad measurements often lack this clinical utility. 

\xhdr{Bias Amplification and Inequities} Inaccurate measurement can amplify bias, reinforcing disparities and inequities in mental health care.

\xhdr{Missed Opportunities for Scientific Progress} Bad measurements limit scientific progress, preventing meaningful contributions and advancements in understanding mental health.

\section{The Path Forward}

By addressing the limitations of current measurement practices in computational mental health research, we hope to create more accurate, robust, and impactful models. To strengthen the scientific rigor and relevance of research, we offer the following recommendations:

\xhdr{Consult Experts} Computational researchers can refer to established guidelines and evidence-based recommendations for the assessment of specific constructs or disorders (e.g., \citet{klein2005toward, antony2005evidence, shear1994standardized}). Collaborating with clinical science colleagues across departments can help guide appropriate measure selection. For projects aiming at immediate clinical application, working closely with a clinician is essential. Clinicians can offer expertise on the tool's clinical utility, including its relevance to real-world practice, ease of use in clinical settings, and alignment with existing diagnostic and treatment workflows. They can also provide feedback on whether the tool offers actionable insights for patient care, supports case conceptualization and treatment planning, and meets the practical needs of diverse clinical populations.


\xhdr{Improve Methodological Rigor When Using Proxies} As previously described, proxy diagnostic signals, such as self-identified diagnoses on social media, have poor validity \cite{ernala2019methodological}. Suggestions for improving the methodological rigor of research using proxies include paring proxy diagnostic signals with offline clinical datasets \cite{inkster2016decade} -- a strong correlation between the proxy and an established clinical outcome, even in just a subsample of participants, could serve to demonstrate the validity of the proxy variable --  and combining multiple proxies to improve reliability \cite{ernala2019methodological}. At the very least, researchers using proxies should clearly state their limitations, e.g., a Twitter-based depression variable should be distinguished from a clinically validated diagnosis, with a note that future research using higher validity measures is needed. 

\xhdr{Adopt Dimensional Measurement} Avoid measuring mental health constructs into binary categories (e.g., ``depressed'' vs. ``not depressed''). Choose dimensional measures that capture severity gradients and avoid dichotomizing continuous variables to form diagnostic categories. Researchers interested in diagnostic status could test this variable in secondary analyses (e.g., \cite{stade2023depression}).

\xhdr{Critically Evaluate Disorder-Specific Measures} Before selecting a construct of interest and its corresponding measure, carefully evaluate whether a disorder-specific approach is necessary. For example, many researchers express interest in indexing anxiety, yet do so using the GAD-7 \cite{spitzer2006brief}, which measures the symptoms of generalized anxiety disorder, a disorder of frequent and uncontrollable worry. Yet uncontrollable worry only represents one form of anxiety pathology. Measures that are not disorder-specific (e.g., MASQ Anxious Symptoms subscale \cite{watson1995testing}) better capture features and processes that cut across the anxiety disorders.  Researchers should use syndrome-specific measures only when this truly aligns with their research goals.

\xhdr{Adopt a Process-Oriented Approach} Instead of focusing solely on specific disorders or syndromes, consider examining transdiagnostic processes. For example, studying constructs that encompass multiple diagnostic categories, such as ``internalizing psychopathology'' or ``fear," can offer more generalizable and integrative insights than research limited to a single diagnosis (e.g., major depression, specific phobia). 


\xhdr{Think Beyond Psychopathology} There are many non  \textit{DSM-5} constructs that are important for health and well-being, especially those that confer risk or protection for psychopathology, such as neuroticism, perfectionism, resilience, and disinhibition. The HiTOP ``components/traits'' level of analysis (e.g., \citet{deyoung2022distinction} offers a starting place for exploring non \textit{DSM-5} constructs.

\xhdr{Maximize Range on Variables of Interest} Since many psychopathology constructs are dimensional, researchers should recruit participants with varying levels of the construct. For example, when studying depression, aim to include the widest possible range of depression severity scores, including subthreshold presentations (e.g., individuals who have symptoms of depression that do not meet \textit{DSM-5} major depressive disorder criteria). Maximizing range on variables of interest should also yield greater effect sizes, since range restriction attenuates effects \cite{linn1968range}.

\xhdr{Attend to Reliability} Reliability sets the upper limit of validity and is crucial for research. Assess reliability using metrics like Cronbach's alpha for self-report \cite{cronbach1951coefficient}, the intraclass correlation coefficient, for dimensional observer ratings, or Cohen's Kappa, for categorical observer ratings \cite{hallgren2012computing}. Training raters thoroughly and enhancing rater competency can ensure good reliability \cite{reichelt2003impact, creed2016beyond}.

\xhdr{Consider Condition Heterogeneity} Mental health conditions are highly heterogeneous, with significant variability in symptom presentation and individual experiences. Two patients with a \textit{DSM-5} diagnosis of major depressive disorder may not share a single symptom \cite{fried2017dep:heterogeneity}. Study designs should account for this variability -- including by analyzing individual symptoms \cite{fried2015depression} -- to avoid oversimplification.

\xhdr{Address Comorbidity} Mental health conditions often co-occur and share overlapping symptoms, hinting that effects thought to be driven by one disorder could be driven by co-occurring conditions. Researchers can account for comorbidity using statistical controls (e.g., \cite{stade2023depression}) or measures that disentangle overlapping conditions (e.g., \cite{watson1995testing}), increasing confidence that effects are unique to a given condition. 

\xhdr{Adopt Longitudinal Measurement} Psychopathology dynamically evolves over time, both in terms of severity and diagnostic label. Longitudinal methods of data collection, and analyses using temporally-aware models or time-series analyses, could help address this reality. Although not historically accounted for in computational research, recent work has begun to examine the relationship between symptoms and language features over time (e.g., \citet{nook2022linguistic}).


We acknowledge that some of the proposed suggestions are challenging for NLP researchers to implement: Computational researcher may require sample sizes prohibitively large for conducting semi-structured diagnostic interviewing (prohibitive from a resource perspective, and even the interaction required to collect self-reported scores, as opposed to social-media based proxy measurement, which may require no interaction between researchers and participants, may be more resource-intensive or involved than is actually feasible. Especially if doing social-media based research requires no interaction with participants whatsoever. Therefore, while advocating for clinician-assessed, dimensional psychopathology measurement as the gold-standard, we suggest that researchers seeking to strengthen measurement approaches can adopt an ``n+1'' approach, where they seek to take one step towards improved measurement. For example, researchers planning to administer a single-item measure can weigh the pros and cons of this approach \cite{allen2022single} and select a measure with demonstrated validity in their population of interest (e.g., \citet{joiner2025reliability}) rather than writing their own item from scratch. Researchers can follow the guidelines for selecting measures in line with transdiagnostic frameworks (e.g. \citet{stanton2020transdiagnostic}) rather than using disorder-specific measures. To demonstrate what different measurement strategies can look like, we present in Table \ref{tab:anxiety-measures} a matrix demonstrating measures of social anxiety that systematically vary on the categories we have highlighted in this paper (validated vs. unvalidated, categorical vs. dimensional, and disorder-specific vs. transdiagnostic). Even incremental improvements can significantly improve validity and utility.

Beyond this, the field sorely needs large, publicly available datasets that include natural language from well-characterized clinical samples, perhaps created leveraging something like a practice-research network \cite{parry2010practice}. Given that many academic mental health clinics routinely administer the same semi-structured interviews, the aggregation of such recordings could be utilized. Diagnostic interviews can be a particularly efficient source of data, because they yield language as well as measurement of psychopathology constructs, and they are often audio recorded. It is possible to use to predict diagnostic severity scores obtained from a separate section of the interview (e.g., \citet{stade2023depression}. 

Accruing this type of large, shared dataset is not without challenges, one of which is the issue of confidentiality. It is difficult to acquire natural language data that are not identifiable or semi-identifiable in some way; and to conduct computational research, this dataset would need clinical variables; risking the disclosure of PHI. However, a potential workaround is not making the raw language public but instead extracting a range of linguistic features (including basic, dictionary-based features as well as more sophisticated, transformer/embedding based features) available.

\section{Discussion and Conclusions}

We highlight challenges in measuring mental health constructs in computational research and propose ways to improve validity. Key issues include overreliance on categorical frameworks, neglect of condition heterogeneity, and inadequate transdiagnostic measures.

Categorical frameworks like the \textit{DSM-5} oversimplify constructs, while dimensional approaches—capturing severity and shared symptoms—enhance model accuracy. Focusing on transdiagnostic constructs, like ``negative affect,'' provides a holistic understanding of mental health.

Condition heterogeneity complicates analysis, but transdiagnostic approaches can address comorbidity and overlapping symptoms. Poor measurement practices introduce errors and biases, so researchers should prioritize validated instruments and diverse datasets.


We argue that researchers should adopt dimensional measures, assess disorder-specific metrics critically, and ensure sample diversity. Implementing transdiagnostic approaches, rater calibration, and reliability checks will further enhance validity.

In conclusion, improving measurement practices is crucial for advancing computational models and mental health care, capturing the complexity of psychopathology, and driving progress in the field.

\newpage
\section{Limitations}


While this work highlights critical issues in the measurement of mental health constructs and provides practical recommendations, it is important to acknowledge its limitations. 

First, although we emphasize the importance of validated measures, we recognize that resource constraints and practical barriers may prevent many researchers from employing clinician-rated assessments or developing fully validated instruments. These barriers underscore the need for scalable, cost-effective alternatives that balance feasibility and validity.

Second, while we center our discussion on practices and pitfalls in the computational research community, we do not mean to imply that computational researchers are the only ones making these mistakes. These measurement issues we outline here are common in behavioral health research more broadly, including psychiatric and psychological research. Addressing these challenges comprehensively will require a broader interdisciplinary effort that includes collaboration across fields.

Third, while we focus on dimensional and transdiagnostic measurement approaches, we acknowledge that these may not be universally applicable. Certain clinical scenarios may necessitate categorical diagnoses for treatment decisions, and some researchers may have justifiable reasons for focusing on specific disorders. Future work should aim to provide clearer guidance on when categorical, dimensional, or transdiagnostic approaches are best.

Fourth, this work does not claim to exhaustively address all the challenges in the measurement of mental health constructs. Other significant issues, such as the influence of cultural biases, ethical considerations in mental health data collection, and the challenges of interpreting results from large-scale datasets, also warrant attention but fall outside the scope of this discussion. 

Finally, while we provide practical recommendations, the field still lacks consensus on "best practices" for measuring mental health constructs in computational research. More empirical studies are needed to evaluate the relative merits of different measurement approaches and their impacts on model performance and real-world applications.

Despite these limitations, we hope that this work stimulates critical reflection and contributes to advancing the validity and utility of mental health research in computational contexts.

\section{Acknowledgements}
Dan Jurafsky  provided critical feedback on an earlier version of this manuscript, and a conversation with Ayelet M. Ruscio shaped our thinking about the different measurement strategies appearing in Table \ref{tab:anxiety-measures}.
This work was supported in part by the Koret Foundation grant for Smart Cities and Digital Living.

\bibliography{custom}

\newpage

\section{Appendix}
\label{sec:appendix}

\begin{table*}[htbp]
\centering
\rotatebox{90}{
\begin{tabular}{|p{2.3cm}|p{4.1cm}|p{4.1cm}|p{4.1cm}|p{4.1cm}|}
\hline
& \multicolumn{2}{c|}{\textbf{Not dimensional}} & \multicolumn{2}{c|}{\textbf{Dimensional}} \\
\hline
& \textbf{Syndrome-specific} & \textbf{Transdiagnostic} & \textbf{Syndrome-specific} & \textbf{Transdiagnostic} \\
\hline
\textbf{Not validated} & r/SocialAnxiety membership status (member vs. not member) \newline\newline Twitter-disclosed social anxiety disorder diagnosis (present vs. absent) \newline\newline Self-reported history of social anxiety disorder diagnosis & r/Anxiety membership status (member vs. not member) \newline\newline Twitter-disclosed anxiety (present vs. absent) & Self-reported social anxiety disorder severity (e.g., ``How much have you been bothered by social anxiety disorder symptoms in the past month, on a 0-10 scale?'') & Self-reported anxiety severity (e.g., ``How much have you been bothered by anxiety symptoms in the past month, on a 0-10 scale?'') \\
\hline
\textbf{Validated} & Clinician-rated SCID-5 social anxiety disorder status (present vs. absent) \newline\newline  Self-reported SPIN score, converted to dichotomous variable reflecting the presence (score > 19) or absence (score $\leq$ 18) of social anxiety disorder & Self-reported MASQ Anxious Symptoms subscale, dichotomized (``high anxious symptoms'' vs. ``low anxious symptoms'') \newline\newline Clinician-rated IMAS fear subfactor score, converted to dichotomous variable reflecting high fear or low fear & Clinician-rated ADIS-5L social anxiety disorder severity on 0 (\textit{none}) to 8 (\textit{very severe}) scale. \newline\newline Self-reported SPIN (0-68 score) \newline\newline Self-reported IDAS-II Social Anxiety Scale & Self-reported MASQ Anxious Symptoms subscale \newline\newline Clinician-rated IMAS fear subfactor \\
\hline
\end{tabular}
}
\caption{
ADIS-5L = Anxiety and Related Disorders Interview
Schedule for \textit{DSM-5} \cite{brown_anxiety_2014}; 
IDAS-II = Inventory of Depression and Anxiety Symptoms \cite{watson2012development}; 
IMAS = Interview for Mood and Anxiety Symptoms \cite{Kotov_Perlman_Gámez_Watson_2015}; 
MASQ = Mood and Anxiety Symptom Questionnaire \cite{watson1995testing}; 
SCID-5 = Structured Clinical Interview for \textit{DSM-5} \cite{first_scid_2016}; 
SPIN = Social Phobia Inventory \cite{connor2000psychometric}.} 

\label{tab:anxiety-measures}
\end{table*}

\end{document}